# DiTFastAttn: Attention Compression for Diffusion Transformer Models

### v1

@@ Abstract
Diffusion Transformers (DiT) excel at image and video generation but face computational challenges due to self-attention’s quadratic complexity. We propose DiTFastAttn, a novel post-training compression method to alleviate DiT’s computational bottleneck. We identify three key redundancies in the attention computation during DiT inference: 1. spatial redundancy, where many attention heads focus on local information; 2. temporal redundancy, with high similarity between neighboring steps’ attention outputs; 3. conditional redundancy, where conditional and unconditional inferences exhibit significant similarity. To tackle these redundancies, we propose three techniques: 1. Window Attention with Residual Caching to reduce spatial redundancy; 2. Temporal Similarity Reduction to exploit the similarity between steps; 3. Conditional Redundancy Elimination to skip redundant computations during conditional generation. To demonstrate the effectiveness of DiTFastAttn, we apply it to DiT, PixArt-Sigma for image generation tasks, and OpenSora for video generation tasks. Evaluation results show that for image generation, our method reduce up to 88% of the FLOPs and achieves up to 1.6x speedup at high resolution generation.
1Tsinghua University    2Infinigence AI    3Carnegie Mellon University    4Shanghai Jiao Tong University
@@ 1 Introduction
Recently, diffusion transformers (DiT) have gained increasing popularity in image (Peebles & Xie, 2023; Chen et al., 2024) and video generation (Open-Sora, 2024). However, a major challenge with DiTs is their substantial computational demand, particularly noticeable when generating high-resolution content. On the one hand, traditional transformer architectures, with their self-attention mechanism, have an $\mathcal{O}(L^{2})$ complexity to the input token length $L$. This quadratic complexity leads to significant computational costs as the resolution of images and videos escalates. As demonstrated in Fig.1, the attention computation becomes the primary computational bottleneck during the inference process as image resolution increases. Specifically, if a $2K\times 2K$ image is tokenized into 16k tokens (Chen et al., 2024), this requires several seconds for attention computation, even on high-end GPUs such as the Nvidia A100. On the other hand, the inference process of diffusion requires a substantial number of neural network inferences due to the multiple denoising steps and the classifier-free guidance (CFG) technique (Ho & Salimans, 2022).
@@
(3) Similarity between the conditional and unconditional inference in attention outputs. We observe that in CFG, the attention outputs of conditional and unconditional inference exhibit significant similarity (SSIM $\geq 0.95$) for certain heads and timesteps. Therefore, we propose the Attention Sharing across CFG (ASC) technique to skip redundant computation during conditional generation.
@@
We conduct extensive experiments to evaluate DiTFastAttn using multiple DiT models, including DiT-XL (Peebles & Xie, 2023) and PixArt-Sigma (Chen et al., 2024) for image generation, and Open-Sora (Open-Sora, 2024) for video generation. Our findings demonstrate that DiTFastAttn consistently reduces the computational cost. Notably, the higher the resolution, the greater the savings in computation and latency. For instance, with PixArt-Sigma, DiTFastAttn delivers a 36% to 88% reduction in attention computation and a speedup of up to 1.6x during the generation of 2048$\times$2048 images.
@@ 2.1 Diffusion Models
Diffusion models (Ho et al., 2020; Rombach et al., 2022; Peebles & Xie, 2023; Chen et al., 2024; Open-Sora, 2024) have gained significant attention due to their superior generative performance compared to GANs (Creswell et al., 2018). Early diffusion models (Ho et al., 2020; Rombach et al., 2022) are implemented based on the U-Net architecture. To achieve better scalability, DiT (Peebles & Xie, 2023) utilizes the transformer architecture instead of U-Net. Diffusion transformer is applied in the fields of image and video generation. PixArt-Sigma (Chen et al., 2024) demonstrates the diffusion transformer’s capability to generate high-resolution images up to 4K. OpenSora (Open-Sora, 2024) presents the diffusion transformer’s ability to generate videos.
@@ 2.3 Local Attention
Various studies have delved into the utilization of local attention patterns, where each token attends to a set of neighboring tokens within a fixed window size, aiming to mitigate the computational burden associated with processing long sequences. The concept of local windowed attention was initially introduced by  Beltagy et al. in Longformer, presenting an attention mechanism that scales linearly with sequence length. Bigbird (Zaheer et al., 2020) extends this idea by incorporating window attention, random attention, and global attention mechanisms, enabling the retention of long-range dependencies while mitigating computational costs. In the realm of computer vision, Swin Transformer (Liu et al., 2021) adopts a similar approach by confining attention computation to non-overlapping local windows, utilizing shifted windows across different layers to capture global context efficiently. Twins Transformer(Chu et al., 2021) and FasterViT(Vasu et al., 2023) employ window-based attention to enhance computational efficiency, leveraging different module designs such as global sub-sampled attention and hierarchical attention to exploit global context effectively. In our work, we employ fixed-sized window attention to accelerate pretrained Diffusion Transformer models and introduce a novel technique named Window Attention with Residual Sharing to preserve long-range dependencies for image tokens.
@@ 3.2 Window Attention with Residual Sharing (WA-RS)
Figure 3: Window Attention with Residual Sharing. (a) Left: Example of the attention map showing the window pattern. Right: The MSE between the window attention outputs in the previous and current step versus the MSE between the output residuals of window and full attention in the previous and current step. The output residual exhibits very small changes over the steps. (b) Computation of Window Attention with Residual Sharing. Window attention that illustrates significant changes is recalculated. Residuals that change minimally are cached and reused in subsequent steps.
@@ 4.1 Settings
We evaluate DiTFastAttn on three commonly used diffusion transformers: DiT (Peebles & Xie, 2023) and Pixart-Sigma (Chen et al., 2024) for image generation tasks, and Open-Sora (Open-Sora, 2024) for video generation tasks. To demonstrate compatibility with fast sampling methods, we build our method upon 20-step DPM-Solver for DiT and Pixart-Sigma, and 100-step IDDPM (Nichol & Dhariwal, 2021) for Open-Sora.
We use ImageNet as the evaluation dataset for calculating quality metrics, and MS-COCO 2017 (Lin et al., 2014) caption as text prompt for Pixart-Sigma models’ image generation. For ImageNet, we generate 5k images to evaluate generation quality. Following previous studies, we employ FID (Heusel et al., 2017), IS (Salimans et al., 2016) and CLIP score (Hessel et al., 2021) as the evaluation metrics. We measure the latency per sample on a single Nvidia A100 40G GPU.
@@ 4.2 Results on Image Generation
Results of Evaluation Metrics and #FLOPs. DiTFastAttn was applied to the pre-trained DiT-2-XL-512, PixArt-Sigma-1024, and PixArt-Sigma-2K models. Figure 5 displays the evaluation results of these models on the ImageNet-5k validation dataset. For the DiT-2-XL-512 and PixArt-Sigma-1024 models, configurations D1, D2, and D3 nearly matched the performance of the original models in terms of IS and FID metrics. The IS and CLIP scores for settings D4, D5, and D6 decrease slightly as a trade-off for achieving higher compression rates. The attention computation of DiT-2-XL-512 was reduced to 88.82%, 76.65%, 67.28%, 58.77%, 50.82%, and 44.28% from D1 to D6, respectively; for PixArt-Sigma-1024, it was reduced to 90.49%, 74.37%, 62.50%, 50.67%, 42.03%, and 36.79%. The PixArt-Sigma-2K model maintained performance close to the original up to D5, with a noticeable performance decline only at D6, where attention computation was reduced to 88.01%, 70.98%, 56.65%, 47.03%, 39.60%, and 35.54%. The comparison of compression effects and evaluation metrics among the three models reveals that as image resolution increases, DiTFastAttn not only achieves greater compression but also better preserves the generative performance of the models.
Visualization of DiTFastAttn’s Generation Results. Figure 6 shows the image generation samples of DiTFastAttn. For the DiT-2-XL-512 and PixArt-Sigma-1024 models, D1, D2, and D3 show visual generation quality similar to that of the original models. While D4, D5, and D6 achieve greater compression and exhibit slight differences in the details of the generated images, they still produce usable quality images. The PixArt-Sigma-2K model maintains image quality comparable to the original up to D4, and configurations D5 and D6 continue to generate high-quality images. This indicates that our compression method, even when reducing attention computation by more than half and compressing to 36% at higher resolutions, still maintains good generation quality.
Figure 5: Image generation performance of DiTFastAttn at various image resolutions under various compression ratios.
@@ 4.3 Results on Video Generation
We apply DitFastAttn on OpenSora for video generation. The results are shown in Figure 10. Specifically, the reduction in computation for configurations labeled D1 through D6 are as follows: 4.29%, 11.51%, 19.55%, 26.74%, 32.62%, 40.37%, respectively. For an extended analysis, additional results are provided in the Appendix.
Table 1: FLOPs fraction and latency fraction of DitFastAttn in Diffusion Transformers comparing with original attention. The latency is evaluated on the Nvidia A100 40G GPU.
@@
DiT-XL-512 1024 FLOPs 50% 69% 35% 0%
Latency 56% 93% 54% 4%
PixArt-Sigma-1024 4096 FLOPs 50% 36% 18% 0%
Latency 52% 73% 40% 2%
PixArt-Sigma-2K 16384 FLOPs 50% 23% 12% 0%
Latency 52% 70% 36% 1%
@@
Compression Results of DiTFastAttn on Various Sequence Lengths. We implement DiTFastAttn based on FlashAttention-2 (Dao, 2023). Table 1 shows the FLOPs fraction and latency fraction of DiTFastAttn in Diffusion Transformers compared with original attention mechanisms. The ASC technique reduces attention computation by 50%, with latency reduction slightly increasing as resolution increases. As resolution increases, WA-RS can reduce attention computation from 69% to 23%, and latency reduction ranges from 93% to 70%. The WA-RS and ASC techniques are orthogonal; they can be used simultaneously without additional overhead.
Overall Latency of DiTFastAttn. Figure 8 shows the latency for image generation and attention as computation decreases when DiTFastAttn is applied. On the DiT-XL-512 model, there is a slight increase in overall latency for both image generation and attention computation in low compression ratio settings. DiTFastAttn introduces no overhead during inference. The latency increase is due to our kernel not being well-implemented, resulting in a slightly lower performance compared to FlashAttention-2. In high compression ratio settings, latency decreases compared to raw, with D6 achieving the lowest latency: overall generation latency is approximately 94.27% of raw, and overall attention latency is 82.32% of raw. For PixArt-Sigma-1024, latency continues to decrease as FLOPs decreases. PixArt-Sigma-2K shows the best performance, with overall generation latency at D6 being 76.32% of raw and overall attention latency at 64.27%. As resolution increases, DiTFastAttn achieves better performance in reducing latency for both overall attention and image generation. It is worth noting that further optimization of our kernel implementation could lead to better latency reductions.
Figure 8: Latency of different resolutions of image generation under different compression ratios.
@@ 5 Conclusion
Limitations and Future Work. First, our method is a post-training compression technique and therefore cannot take advantage of training to avoid the performance drop. Second, our greedy compression method selection is simple but may not find the optimal. Third, our method only reduces the cost of attention module. In future work, we plan to explore training-aware compression methods. We also aim to extend our approach to other modules. Additionally, further kernel-level optimizations may unlock even greater speedups for the proposed compression techniques.
@@ Acknowledgements
This work was supported by National Natural Science Foundation of China (No. 62325405, 62104128, U19B2019, U21B2031, 61832007, 62204164), Tsinghua EE Xilinx AI Research Fund, and Beijing National Research Center for Information Science and Technology (BNRist). We thank for all the support from Infinigence-AI.
@@ A.1 Societal Impacts
A.2 Release Safeguards
While DiTFastAttention itself does not release new pre-trained models, the compression capabilities it provides could enable easier sharing and deployment of powerful diffusion models that have risks of misuse. To mitigate risks of misuse, we have implemented access control. Users must agree to terms prohibiting unethical applications.
A.3 Codes
Our codebase is anonymously available in the Supplementary Materials. For implementation details, we have them added to readme documentation in the codebase.
A.4 Details of method for deciding the compression plan
We employ a greedy algorithm to determine the appropriate compression strategy for each transformer layer and denoising step. For each layer and step, we evaluate the error between the model outputs with and without compression using each of the four candidate strategies: Attention Sharing across Timesteps (AST), Window Attention with Residual Sharing (WA-RS) and Attention Sharing across CFG (ASC), WA-RS alone, and ASC alone. For each timestep and layer, we select the strategy that achieves the highest computation reduction while keeping the loss below a predefined threshold $\frac{i}{|M|}\delta$.
[…]

### v2

@@ Abstract
Diffusion Transformers (DiT) excel at image and video generation but face computational challenges due to the quadratic complexity of self-attention operators. We propose DiTFastAttn, a post-training compression method to alleviate the computational bottleneck of DiT. We identify three key redundancies in the attention computation during DiT inference: (1) spatial redundancy, where many attention heads focus on local information; (2) temporal redundancy, with high similarity between the attention outputs of neighboring steps; (3) conditional redundancy, where conditional and unconditional inferences exhibit significant similarity. We propose three techniques to reduce these redundancies: (1) Window Attention with Residual Sharing to reduce spatial redundancy; (2) Attention Sharing across Timesteps to exploit the similarity between steps; (3) Attention Sharing across CFG to skip redundant computations during conditional generation. We apply DiTFastAttn to DiT, PixArt-Sigma for image generation tasks, and OpenSora for video generation tasks. Our results show that for image generation, our method reduces up to 76% of the attention FLOPs and achieves up to 1.8$\times$ end-to-end speedup at high-resolution (2k $\times$ 2k) generation.
1Tsinghua University    2Infinigence AI    3Shanghai Jiao Tong University
@@ 1 Introduction
Recently, diffusion transformers (DiT) have gained increasing popularity in image (Peebles & Xie, 2023; Chen et al., 2024) and video generation (Brooks et al., 2024). However, a major challenge with DiTs is their substantial computational demand, particularly noticeable when generating high-resolution content. On the one hand, traditional transformer architectures, with their self-attention mechanism, have an $\mathcal{O}(L^{2})$ complexity to the input token length $L$. This quadratic complexity leads to significant computational costs as the resolution of images and videos escalates. As demonstrated in Fig.1, the attention computation becomes the primary computational bottleneck during the inference process as image resolution increases. Specifically, if a $2K\times 2K$ image is tokenized into 16k tokens (Chen et al., 2024), this requires several seconds for attention computation, even on high-end GPUs such as the Nvidia A100. On the other hand, the inference process of diffusion requires a substantial number of neural network inferences due to the multiple denoising steps and the classifier-free guidance (CFG) technique (Ho & Salimans, 2022).
@@
(3) Similarity between the conditional and unconditional inference in attention outputs. We observe that in CFG, the attention outputs of conditional and unconditional inference exhibit significant similarity (SSIM $\geq 0.95$) for certain heads and timesteps. Therefore, we propose the Attention Sharing across CFG (ASC) technique to skip redundant computation during unconditional generation.
@@
We conduct extensive experiments to evaluate DiTFastAttn using multiple DiT models, including DiT-XL (Peebles & Xie, 2023) and PixArt-Sigma (Chen et al., 2024) for image generation, and Open-Sora (Open-Sora, 2024) for video generation. Our findings demonstrate that DiTFastAttn consistently reduces the computational cost. Notably, the higher the resolution, the greater the savings in computation and latency. For instance, with PixArt-Sigma, DiTFastAttn delivers a 20% to 76% reduction in attention computation and a end-to-end speedup of up to 1.8x during the generation of 2048$\times$2048 images.
@@ 2.1 Diffusion Models
Diffusion models (Ho et al., 2020; Rombach et al., 2022; Peebles & Xie, 2023; Chen et al., 2024; Brooks et al., 2024) have gained significant attention due to their superior generative performance compared to GANs (Creswell et al., 2018). Early diffusion models (Ho et al., 2020; Rombach et al., 2022) are implemented based on the U-Net architecture. To achieve better scalability, DiT (Peebles & Xie, 2023) utilizes the transformer architecture instead of U-Net. Diffusion transformer is applied in the fields of image and video generation. PixArt-Sigma (Chen et al., 2024) demonstrates the diffusion transformer’s capability to generate high-resolution images up to 4K. Sora (Brooks et al., 2024) presents the diffusion transformer’s ability to generate videos.
@@ 2.3 Local Attention
Various studies have delved into the utilization of local attention patterns, where each token attends to a set of neighboring tokens within a fixed window size, aiming to mitigate the computational burden associated with processing long sequences. The concept of local windowed attention was initially introduced by  Beltagy et al. in Longformer, presenting an attention mechanism that scales linearly with sequence length. Bigbird (Zaheer et al., 2020) extends this idea by incorporating window attention, random attention, and global attention mechanisms, enabling the retention of long-range dependencies while mitigating computational costs. In the realm of computer vision, Swin Transformer (Liu et al., 2021) adopts a similar approach by confining attention computation to non-overlapping local windows, utilizing shifted windows across different layers to capture global context efficiently. Twins Transformer(Chu et al., 2021), FasterViT(Vasu et al., 2023), and Neighborhood attention transformer (Hassani et al., 2023) employ window-based attention to enhance computational efficiency, leveraging different module designs such as global sub-sampled attention and hierarchical attention to exploit global context effectively. In our work, we employ fixed-sized window attention to accelerate pretrained Diffusion Transformer models and introduce a novel technique named Window Attention with Residual Sharing to preserve long-range dependencies for image tokens.
@@ 2.4 Attention Sharing
2.5 Other Methods to Accelerate Diffusion Models
Network quantization is a widely used technique for reducing the bitwidth of weights and activations, effectively compressing both image generation models (Shang et al., 2023; Zhao et al., 2024b) and video generation models (Zhao et al., 2024a). Scheduler optimization is another popular approach aimed at decreasing the number of timesteps in the denoising process (Song et al., 2020; Lu et al., 2022; Liu et al., 2023a). Additionally, distillation serves as an effective method for minimizing the timesteps required during denoising (Salimans & Ho, 2022; Meng et al., 2023; Liu et al., 2023b). DiTFastAttn offers a complementary solution, as it operates independently of the specific quantization bitwidth, scheduler, and timesteps employed.
@@ 3.2 Window Attention with Residual Sharing (WA-RS)
Figure 3: Window Attention with Residual Sharing. (a) Left: Example of the attention map showing the window pattern. Right: The MSE between the window attention outputs in the previous and current step (yellow line) versus the MSE between the output residuals of window and full attention in the previous and current step (blue line). The output residual exhibits minimal changes over the steps. (b) Computation of Window Attention with Residual Sharing. Window attention that illustrates significant changes is recalculated. Residuals that change minimally are cached and reused in subsequent steps.
@@ 4.1 Settings
We evaluate DiTFastAttn on three commonly used diffusion transformers: DiT (Peebles & Xie, 2023) and Pixart-Sigma (Chen et al., 2024) for image generation tasks, and Open-Sora (Open-Sora, 2024) for video generation tasks. To demonstrate compatibility with fast sampling methods, we build our method upon 50-step DPM-Solver for DiT and Pixart-Sigma, and 200-step IDDPM (Nichol & Dhariwal, 2021) for Open-Sora.
For calculating quality metrics, we use ImageNet as the evaluation dataset for DiT and MS-COCO as the evaluation dataset for PixArt-Sigma. MS-COCO 2014 caption is used as text prompt for Pixart-Sigma models’ image generation. To evaluate generation quality, we generate 50k images for DiT models and 30k images for PixArt-Sigma models. Following previous studies, we employ FID (Heusel et al., 2017), IS (Salimans et al., 2016) and CLIP score (Hessel et al., 2021) as the evaluation metrics. We measure the latency per sample on a single Nvidia A100 GPU.
@@ 4.2 Results on Image Generation
Results of Evaluation Metrics and Attention FLOPs. DiTFastAttn was applied to the pre-trained DiT-XL-2-512, PixArt-Sigma-1024, and PixArt-Sigma-2K models. Table 1 displays the evaluation results of these models. For the DiT-XL-2-512 and PixArt-Sigma-1024 models, configurations D1, D2, and D3 nearly matched the performance of the original models in terms of IS and FID metrics. Comparison of compression effects and evaluation metrics between the three models reveals that as image resolution increases, DiTFastAttn not only achieves greater compression but also better preserves the generative performance of the models.
Compression Plan after Search. Figure 5 illustrates the compression plan under the D6 setting. For the DiT model, AST and ASC are utilized in the early timesteps, with full attention primarily appearing in the initial attention layers. In contrast, the PixArt-Sigma model employs AST sporadically in the first two layers and in the middle attention layers during the intermediate timesteps, while the combination of WA-RS and ASC is notably predominant in the final timesteps. This variability in the distribution of different types of redundancy across models highlights the absence of a universal compression strategy, underscoring the necessity for tailored plan searches. Additional compression plans for other settings are provided in the Appendix.
Figure 5: Compression plan for DiT-XL-512, PixArt-Sigma-XL-1024 and PixArt-Sigma-XL-2K at D6 with the number of DPM-Solver steps set to 50.
Visualization of DiTFastAttn’s Generation Results. Figure 6 presents image generation samples from DiTFastAttn. The D1, D2, and D3 configurations of the DiT-XL-2-512 and PixArt-Sigma-1024 models demonstrate visual generation quality comparable to the original models. In contrast, D4, D5, and D6 achieve greater compression, exhibiting slight variations in detail while still producing acceptable-quality images. The PixArt-Sigma-2K model maintains image quality similar to the original up to D4, with D5 and D6 also generating high-quality outputs. This suggests that our compression method effectively preserves generation quality, even when reducing attention computation by over 50% and compressing to 33% at higher resolutions.
Table 1: Image generation performance of DiTFastAttn at various image resolutions under various compression ratios. The FID, IS, and CLIP results are marked in different makers. The "Attn FLOPs" represents the fraction of computation in the multi-head attention module compared to the raw model.
Model DiT-XL-2 512x512 PixArt-Sigma-XL 1024x1024 PixArt-Sigma-XL 2048x2048
Score IS FID
Attn
FLOPs
IS FID CLIP
Attn
FLOPs
IS FID CLIP
Attn
FLOPs
Raw 408.16 25.43 100% 24.33 55.65 31.27 100% 23.67 51.89 31.47 100%
D1 412.24 25.32 85% 24.27 55.73 31.27 90% 23.28 52.34 31.46 81%
D2 412.18 24.67 69% 24.25 55.69 31.26 74% 22.90 53.01 31.32 60%
D3 411.74 23.76 59% 24.16 55.61 31.25 63% 22.96 52.54 31.36 46%
D4 391.80 21.52 49% 24.07 55.32 31.24 52% 22.95 51.74 31.39 36%
D5 370.07 19.32 41% 24.17 54.54 31.22 44% 22.82 51.21 31.34 29%
D6 352.20 16.80 34% 23.94 52.73 31.18 37% 22.38 49.34 31.28 24%
@@ 4.3 Results on Video Generation
We apply DitFastAttn on OpenSora for video generation with five thresholds from 0.01 to 0.05. The results are shown in Figure 7. Specifically, the reduction in computation for configurations labeled D1 through D6 are as follows: 7.63%, 19.50%, 30.16%, 37.66%, 40.52%, respectively. For an extended analysis, additional results are provided in the Appendix.
Table 2: FLOPs fraction and latency fraction of DitFastAttn in Diffusion Transformers comparing with original attention. The latency is evaluated on the Nvidia A100 GPU.
@@
DiT-XL-2 512x512 1024 Attn FLOPs 50% 77% 38% 0%
Attn Latency 59% 85% 51% 4%
PixArt-Sigma-XL 1024x1024 4096 Attn FLOPs 50% 51% 26% 0%
Attn Latency 54% 54% 31% 3%
PixArt-Sigma-XL 2048x2048 16384 Attn FLOPs 50% 33% 16% 0%
Attn Latency 52% 35% 19% 1%
@@
Compression Results of DiTFastAttn on Various Sequence Lengths. We implement DiTFastAttn based on FlashAttention-2 (Dao, 2023). Table 2 shows the FLOPs fraction and latency fraction of DiTFastAttn in Diffusion Transformers compared with original attention mechanisms. The ASC technique reduces attention computation by 50%, with latency reduction slightly increasing as resolution increases. As resolution increases, WA-RS can reduce attention computation from 77% to 33%, and latency reduction ranges from 85% to 35%. The WA-RS and ASC techniques are orthogonal; they can be used simultaneously without additional overhead.
Overall Latency of DiTFastAttn. Figure 8 shows the latency for image generation and attention as computation decreases when DiTFastAttn is applied. DiTFastAttn achieves end-to-end latency reduction for all three model at all compression ratio settings. PixArt-Sigma-2K shows the best performance, with overall generation latency at D6 being 56% of raw and overall attention latency at 37%. The result indicates that as resolution increases, DiTFastAttn achieves better performance in reducing latency for both overall attention and image generation.
Figure 8: Latency of different resolutions of image generation under different compression ratios. DiT runs with a batch size of 8, while PixArt-Sigma models with a batch size of 1. The blue line delineates the latency for end-to-end image generation, whereas the orange line represents the latency of multi-head attention module.
@@ 5 Conclusion
Limitations. First, our method is a post-training compression technique and therefore cannot take advantage of training to avoid the performance drop. Second, our method mainly focuses on inference acceleration instead of VRAM reduction. When AST is applied, the attention hidden states from previous timestep will be stored and will bring extra VRAM usage. Third, our simple compression method may not find the optimal compression plan. Fourth, our method only reduces the cost of attention module.
@@ Acknowledgements
This work was supported by National Natural Science Foundation of China (No. 62325405, 62104128, U19B2019, U21B2031, 61832007, 62204164), Tsinghua EE Xilinx AI Research Fund, and Beijing National Research Center for Information Science and Technology (BNRist).
@@ A.1 Societal Impacts
A.2 Details of the Method for Deciding the Compression Plan